OpenStudy (anonymous):

20.) A side of a square has length 3. Find the length of its diagonal. 3 3 sqrt(2) 1.5 3 sqrt(3)

OpenStudy (dumbcow):

3^2 +3^2 = d^2 find d

OpenStudy (anonymous):

Since each side is 3, you can cut the square in half to form two right triangles, then use the pythagorean theorem. Therefore, \[3^2+3^2=c^2\]which gives you \[c^2=18\]and therefore the length of the diagonal is \[\sqrt(18)\]

OpenStudy (anonymous):

sqrt(18) = 18 = 3 * 3 * 2 groups of 2 items that are alike come out side the radical sign. 3sqrt(2)